# Community detection for Contexual-LSBM: Theoretical limitation on misclassfication ratio and effecient algorithm

### v1

@@ Abstract
The integration of both network information and node attribute information has recently gained significant attention in the context of community recovery problems. In this work, we address the task of determining the optimal classification rate for the Label-SBM(LSBM) model with node attribute information and. Specifically, we derive the optimal lower bound, which is characterized by the Chernoff-Hellinger divergence for a general LSBM network model with Gaussian node attributes. Additionally, we highlight the connection between the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in our model and those introduced in Yun and Proutiere (2016) and Lu and Zhou (2016). We also presents a consistent algorithm based on spectral method for the proposed aggreated latent factor model.
@@ 1 Introduction
The community recovery problem Girvan and Newman (2002); Newman and Leicht (2007); Handcock et al. (2007); Fortunato (2010); Karrer and Newman (2011) is one of the fundamental challenges in the field of network science and data analysis. It involves identifying and reconstructing the underlying community structure within a network or graph. Accurately recovering these communities is crucial for understanding the organization and functionality of complex systems, ranging from social networks and biological systems to communication networks and financial markets.
Among the various proposed models, the Stochastic Block Model (SBM) Holland et al. (1983) stands out as one of the most significant and extensively studied in recent decades. Given known parameters $n$ and $K$, according to the definition of the SBM, each node with index $i\in[n]$ belonging to the set of nodes $\mathcal{V}$ is independently assigned to one of the $[K]$ clusters with probability $\alpha_{k}$. The probability of observing an edge $A_{ij}$ between any two nodes $i,j\in[N]$ is characterized by $\mathbb{P}_{\sigma(i),\sigma(j)}$ for a given label indicator function $\sigma(\cdot):[N]\rightarrow[K]$.
In addition to the basic SBM, numerous variants have been proposed and investigated, such as weighted/valued SBM Jog and Loh (2015); Aicher et al. (2015); Kurihara et al. (2006); Nowicki and Snijders (2001); Xu et al. (2020), degree-corrected SBM Karrer and Newman (2011); Gao et al. (2018); Yan et al. (2014); Qin and Rohe (2013), overlaping/mixed-membership SBM Airoldi et al. (2008); Latouche et al. (2014); Fu et al. (2009); Xing et al. (2010), hypergraph SBM Ghoshdastidar and Dukkipati (2014, 2015, 2017); Zhang and Tan (2023); Cole and Zhu (2020); Nandy and Bhattacharya (2024); Wang (2023); Gaudio and Joshi (2023), latent space model Hoff et al. (2002); Rohe et al. (2011); Smith et al. (2019); Koren (2005); Christensen and Golino (2021); Bing et al. (2020); Rohe and Zeng (2023); Bing et al. (2023) and so on. Among the aformentioned variants, Label-SBM(LSBM) Heimlicher et al. (2012), an extension of the SBM in terms of the non-binary observed edges, has recently garnered significant attention due to its capability in modeling diverse interaction types. More specfically speaking, the observed edge $l$ between any two nodes $u,v$ is sampled from a set ${0,1,2\cdots,L}$, where the probability is conditioned on the clusters $\sigma(u),\sigma(v)$ and and the label $l$, denoted as $\mathbf{P}(\sigma(u),\sigma(v),l)$. Moreover, real-world applications necessitate the analysis of contextual networks, as side information such as node attributes can be instrumental in uncovering the identity of node labels within groups. Substantial research has been dedicated to this area, as seen in Zhang et al. (2016); Abbe et al. (2022); Braun et al. (2022); Yan and Sarkar (2021). However, to the best of our knowledge, rigorous analysis of the integration of both types of information—Label-SBM and node attributes—has not yet been conducted.
In this paper, we investigate a model that combines the Label-Stochastic Block Model (LSBM) with contextual node attributes, referred to as the Contextual-Label-SBM (CLSBM). We assume a Gaussian mixture model for the contextual attribute vectors. Our primary focus is on analyzing the lower bound of the misclassification rate based on the model’s parameters. Specifically, let $s$ denote the expected number of misclassified nodes. Our objective is to establish the lower bound of the misclassification rate for the CLSBM when $s=o(n)$. While this result has been established for the LSBM model in (Yun and Proutiere, 2016, Theorem 1), our Theorem 1 is the first to extend this result to the CLSBM. Additionally, we aggregate topological and contextual information into a latent factor model and apply a spectral-based method to consistently estimate the community assignments for each node. The paper is organized as follows: In Section 2, we review related literature and outline our main contributions. In Section 3.1, we formalize the model definitions and discuss the assumptions necessary for our theoretical findings. Section 3.2 provides the theoretical lower bound for the number of misclassified nodes in the model and explores its connections with previous models. Finally, Section 3.3 introduces a spectral clustering algorithm for the proposed aggregated factor model and presents theoretical guarantees for its performance.
Notations
@@
In this paper, bold lowercase letters, such as $\mathbf{x}$, represent vectors, while bold uppercase letters, like $\mathbf{X}$, denote matrices and $3$-d arrays. For a matrix $\mathbf{X}$, $\mathbf{X}_{ij}$ indicates the element located at the $i$-th row and $j$-th column, with $\mathbf{X}_{i\cdot}$ and $\mathbf{X}_{\cdot j}$ representing the $i$-th row and $j$-th column of $\mathbf{X}$, respectively. Similarly, for a $3$-d array $\mathbf{X}$, we use $\mathbf{X}_{i,j,k}$ to indicate the element at indices $i,j,k$ for each dimension. Occasionally, $\mathbf{x}_{j}$ is used as a simplified notation for $\mathbf{X}_{\cdot j}$.
@@
The operators of probability and expectation are denoted by $\mathbb{P}$ and $\mathbb{E}$, respectively. For any integer $K\in\mathbb{N}$, the set $\{1,2,3,\cdots,K\}$ is represented by $[K]$. The notation $\left\|\mathbf{x}\right\|_{q}$, $\left\|\mathbf{X}\right\|_{F}$, $\left\|\mathbf{X}\right\|_{op}$ represent the $\ell_{p}$ norm of a vector $\mathbf{x}$, Frobenius norm and operoator norm of matrix $\mathbf{X}$ respectively. Here we use $\mathbf{1}(\cdot)$ to display the indicator function. For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}\lesssim b_{n}$ implies $a_{n}\leq Cb_{n}$ for some constant $C$.
2 Related work and main contribution
@@
Jog and Loh (2015) was the first to investigate the threshold of exact recovery for the labeled/weighted SBM model by extending the results from Zhang and Zhou (2016). Additionally, Yun and Proutiere (2016) derived the optimal misclassification proportion for the LSBM and introduced a two-stage algorithm to achieve the information-theoretic limit. Besides, the contextual SBM, a generalization of SBM that incorporates additive node attributes to provide further information for identifying community clusteres, has been explored in various studies Braun et al. (2022); Deshpande et al. (2018); Dreveton et al. (2024); Abbe et al. (2022). However, to the best of our knowledge, the optimal classification rate for models with general non-binary labeled edges combined with independent arbitrary node attributes still remains unexplored. Our Theorem 1 shows that the lower bound for the expected misclasification number of nodes (as defined in Eq. 2) is $n\exp\left(-(1+\epsilon_{n})D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})n\right)$ for some sequence $\epsilon_{n}$ approaching zero. Here the key quantity $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$(defined in (4)) is the divergence defined based on the minimal value for the sum of two KL divergences, representing the topological and attribute information respectively. Furthermore, Remark 1 reveals a connection between the minimal value of sum of KL divergences over constrained sets and the Chernoff divergence, extending the results of (Yun and Proutiere, 2016, Claim $4$) to our settings and consequently prove that the aformentioned lower bound indeed match the optimal classfication rate in (Label)SBM(Zhang and Zhou (2016); Yun and Proutiere (2016)) and Gaussian mixture model(Lu and Zhou (2016)). One of our main technical contribution stems in the analysis for the solve of the problem (4). It is worth noting that, while the optimal solution for $D_{A}$ has been investigated in Yun and Proutiere (2016), the impact of the newly introduced term $D_{X}$ on the main results remains unclear. The additional term introduces further challenges, as it requires solving an optimization problem in function spaces, which is significantly more complex. In addition to establishing the theoretical lower bound, we also apply spectral methods to the proposed aggregated latent factor model, achieving a polynomial error rate. While this result is not optimal, the output of the spectral method can serve as a valuable initialization for further refinement in a second step, e.g see Gao et al. (2018); Yun and Proutiere (2016); Braun et al. (2022).
@@ 3.1 Model definition
We primarily follow the notation introduced in Yun and Proutiere (2016). Each node is independently assigned to a cluster $\mathcal{V}_{k}$ with probability $\alpha_{k}$, where $k\in[K]$, and this assignment is independent of the node number $n$. Each edge $A_{i,j}\in\mathcal{V}^{2}$ is labeled with $l\in\{0,1,\cdots,L\}$, where the labeling probability depends on $i$, $j$, and $l$, and is denoted as $\mathbf{P}(\sigma(i),\sigma(j),l)$. Here, $\mathbf{P}\in\mathbb{R}^{K\times K\times(L+1)}$, and $\sigma(\cdot)$ is the node cluster indicator function. Here we summerize the above in the following definition of LSBM model as follows:
Definition 1 (Label Stochastic Block Model).
@@
Let $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$ be an $3$-d tensor and we write $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$ if $\forall l\in\{0,1,\cdots,L\}$ $\mathbf{A}_{l}\in\left\{0,1\right\}^{n\times n}$ is symmertical and its upper triangular has independent elements such that
$\mathbb{P}(\mathbf{A}_{l}(i,j)=1)=\mathbf{P}(\sigma(i),\sigma(j),l),~{}\sum_{l=0}^{L}\mathbf{A}_{l}(i,j)=1,~{}\textit{and}\quad\mathbb{P}(\sigma(i)=k)=\boldsymbol{\alpha}_{k},$ (1)
$\forall i\neq j\in[n],\,k\in[K]$. Here we assume $\mathbf{A}_{l}(i,i)=0$ for all $i\in[K],l\in[L]$ to remove the self loop of the graph, $\boldsymbol{\alpha}$ is a constant vector independent of $n$ and $\sigma(\cdot):[n]\rightarrow[K]$ is the cluster assignment indicator function.
It is worth mentioning here that both $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$ and any third dimension of $\mathbf{P}$, such as $\mathbf{P}_{ij\cdot}\in\mathbb{R}_{+}^{(L+1)}$, must correspond to a probability distribution.
@@
In addition to topological information, contextual attributes are also incorporated into the final model. We assume that each node $i$ is equipped with an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the network $\mathbf{A}$ is considered independent of the sets of attribute vectors $\{\mathbf{x}_{i}\}$ given the label function $\sigma(\cdot)$. This setting was initially proposed in Deshpande et al. (2018) and has since been discussed in subsequent works Binkiewicz et al. (2017); Yan and Sarkar (2021); Weng and Feng (2016).
Definition 2 (Contexual Label Stochastic Block Model).
@@
Let $d\geq 2$, $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$, $\boldsymbol{\mu}\in\mathbb{R}^{d\times K}$ ,and we write ($\left(\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}},\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\right)\sim CLSBM(L,d,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot))$, if $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$, and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\sim\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
Here we assume $\left\{\mathbf{A}_{l}\right\}_{l\in[L]}$ and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ are independently generated given the parameters respectively. Assume there exist a assignment vector fucntion $\bar{\sigma}(\cdot)$ to indicate the assignment for each node and this $\bar{\sigma}(\cdot)$ is estimated from the observed labels $\left\{\mathbf{A}_{l}\right\}_{l\in[L]\bigcup\{0\}]}$ and attribute $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$. In the next section, we provide the information theoretical lower-bound for the misclassification rate between estimated $\bar{\sigma}(\cdot)$ and ground truth $\sigma(\cdot)$ such that
@@
$\bar{s}:=\inf_{\begin{subarray}{c}\pi\in S_{n}\\ \bar{\sigma}\in\mathcal{F}_{a}\end{subarray}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (2)
@@
Here, $S_{n}$ represents the set of all permutations of $[K]$, and $\mathcal{F}_{a}$ denotes the set of all assignment functions $\sigma:[N]\to[K]$. Before presenting the main results, we first outline our assumptions.
@@
$\forall l\in\{0,1,\cdots,L\},~{}\forall i,j,k\in[K],\quad\frac{\mathbf{P}(i,j,l)}{\mathbf{P}(i,k,l)}\leq\eta_{1},\quad\textit{and},\quad\left\|\boldsymbol{\mu}_{i}\right\|_{2}\leq\eta_{2}.$ (3)
The assumption mentioned earlier arises from the limitations of our proof techniques, and we believe it could potentially be eliminated with improved proof mechanisms. Notably, the assumption regarding $\mathbf{P}$ originates from Yun and Proutiere (2016), while the assumption for $\boldsymbol{\mu}$ is introduced to bound the second-order moments of the pseudolikelihood estimator for the perturbed model in the proof of Theorem 1. It is also worth mentioning that similar assumptions on $\boldsymbol{\mu}$ are commonly found in previous literature, such as Abbe et al. (2022).
3.2 Lower bound on number of misclassfication nodes
In this section we provide our main result. We start with the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$:
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{k_{1}\neq k_{1}\in[K]}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}_{+}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ (4)

### v3

# Community Detection for Contextual-LSBM: Theoretical Limitations of Misclassification Rate and Efficient Algorithms

@@ Abstract
The integration of network information and node attribute information has recently gained significant attention in the community detection literature. In this work, we consider community detection in the Contextual Labeled Stochastic Block Model (CLSBM), where the network follows an LSBM and node attributes follow a Gaussian Mixture Model (GMM). Our primary focus is the misclassification rate, which measures the expected number of nodes misclassified by community detection algorithms. We first establish a lower bound on the optimal misclassification rate that holds for any algorithm. When we specialize our setting to the LSBM (which preserves only network information) or the GMM (which preserves only node attribute information), our lower bound recovers prior results. Moreover, we present an efficient spectral-based algorithm tailored for the CLSBM and derive an upper bound on its misclassification rate. Although the algorithm does not attain the lower bound, it serves as a reliable starting point for designing more accurate community detection algorithms (as many algorithms use spectral method as an initial step, followed by refinement procedures to enhance accuracy).
@@ 1 Introduction
The community detection problem Girvan and Newman (2002); Newman and Leicht (2007); Fortunato (2010); Karrer and Newman (2011) is one of the fundamental challenges in the field of network science and data analysis. It focuses on identifying and reconstructing the underlying community structure within a network or graph. Accurately identifying these communities is crucial for understanding the structure and function of complex systems, ranging from social networks and biological systems to communication networks and financial markets.
Among the various proposed models, the Stochastic Block Model (SBM) Holland et al. (1983) stands out as one of the most significant and extensively studied in recent decades. Given parameters $n$ and $K$, according to the definition of the SBM, each node with index $i\in[n]$ belonging to the set of nodes $\mathcal{V}$ is independently assigned to the $k$-th community (where $k\in[K]$) with probability $\alpha_{k}$. Denote $\sigma:[n]\rightarrow[K]$ as the ground truth indicator function such that $\sigma(i)=k$ if the node $i$ belongs to community $k$, thus the community detection problem is to infer the ground truth indicator function $\sigma(\cdot)$ from the observed graph (usually represented by an adjacency matrix $\mathbf{A}$). Here the probability of observing an edge $\mathbf{A}_{ij}$ between any two nodes $i,j\in[n]$ is characterized by $\mathbb{P}_{\sigma(i),\sigma(j)}$. In addition to the basic SBM, many variants have been proposed and investigated, such as weighted/labeled SBM Jog and Loh (2015); Aicher et al. (2015); Nowicki and Snijders (2001); Yun and Proutiere (2016), degree-corrected SBM Karrer and Newman (2011); Gao et al. (2018); Yan et al. (2014); Qin and Rohe (2013), overlapping/mixed-membership SBM Airoldi et al. (2008); Latouche et al. (2014); Fu et al. (2009); Xing et al. (2010), hypergraph SBM Ghoshdastidar and Dukkipati (2014); Zhang and Tan (2023); Cole and Zhu (2020); Nandy and Bhattacharya (2024), latent space model Hoff et al. (2002); Rohe et al. (2011); Bing et al. (2020); Rohe and Zeng (2023); Bing et al. (2023). Among the aforementioned variants, in this paper we focus on the labeled SBM (LSBM) Heimlicher et al. (2012); Yun and Proutiere (2016), an extension of the SBM in terms of the non-binary observed edges. Compared to the basic SBM, the non-binary edge formulation provides enhanced capability in modeling diverse types of interactions between any pair of nodes.
In the community detection literature, many existing methods rely solely on topological information, such as adjacency or Laplacian matrices. However, real-world applications often require the analysis of contextual networks, as side information like node attributes can play a crucial role in uncovering node labels within communities. Significant research has been dedicated to this area, as evidenced by works such as Zhang et al. (2016); Abbe et al. (2022); Braun et al. (2022); Yan and Sarkar (2021). Specifically, Zhao et al. (2021); Sima et al. (2021); Abbe et al. (2022); Dreveton et al. (2024); Braun et al. (2022) investigate the threshold for exact recovery (equavilent to strong consistency Abbe (2018)) of the Contextual-SBM, with Zhao et al. (2021); Sima et al. (2021); Abbe et al. (2022) focusing on the special case of two communities and Dreveton et al. (2024); Braun et al. (2022) extending the results to the more general case of $K$ communities. While exact recovery is a valid performance metric that has received significant theoretical attention, it requires the number of misclassified nodes to be exactly zero, which is overly stringent and often impractical. In contrast, most real-world applications of community detection or node classification often prioritize weak consistency, where a small fraction of nodes is allowed to be misclassified. This aligns more closely with practical needs, where partial misclassification is acceptable and more realistic.
Motivated by practical considerations, in this paper, we investigate a model that integrates the LSBM with contextual Gaussian node attributes, referred to as Contextual-LSBM (CLSBM). Our focus is on the relationship between the misclassification rate and the model parameters. Specifically, let $s$ represent the expected number of misclassified nodes. Our primary objective is to establish the lower bound of the misclassification rate for the CLSBM when $s=o(n)$. While this result has been established for the LSBM in (Yun and Proutiere, 2016, Theorem 1), our Theorem 1 is the first to extend this result to the setting where contextual node attributes are present. Notably, Braun et al. (2022) and Dreveton et al. (2024) also investigate contextual models similar to ours. However, neither Braun et al. (2022) nor Dreveton et al. (2024) allows edges to carry labels, and Dreveton et al. (2024) focuses specifically on achieving exact recovery. These can be viewed as special cases or reductions of our more general results (see Theorem 1 and Remark 1). In addition to the lower bound, we also design an efficient community detection algorithm for CLSBM. The algorithm begins by aggregating topological and contextual information into a latent factor model and then employs a spectral-based method to consistently estimate the community assignments for each node.
1.1 Organization
The remainder of the paper is organized as follows: In Section 2, we review related literature and highlight our main contributions. In Section 3.1, we formalize the model definitions and discuss the assumptions underpinning for our theoretical findings. Section 3.2 provides the theoretical lower bound for the number of misclassified nodes in the model and explores its connections with earlier work. Furthermore, Section 3.3 introduces an efficient spectral community detection algorithm for the proposed aggregated factor model and presents theoretical guarantees for its performance. Finally we conclude our results and future directions in Section 4.
1.2 Notations
@@
In this paper, bold lowercase letters, such as $\mathbf{x}$, represent vectors, while bold uppercase letters, like $\mathbf{X}$, denote matrices and $3$-dimensional arrays. For a matrix $\mathbf{X}$, $\mathbf{X}_{ij}$ indicates the element located at the $i$-th row and $j$-th column, with $\mathbf{X}_{i\cdot}$ and $\mathbf{X}_{\cdot j}$ representing the $i$-th row and $j$-th column of $\mathbf{X}$, respectively. For a $3$-dimentional array $\mathbf{X}$, we use $\mathbf{X}(i,j,k)$ to indicate the element at indices $i,j,k$ for each dimension. Occasionally, $\mathbf{x}_{j}$ is used as a simplified notation for $\mathbf{X}_{\cdot j}$.
@@
The operators of probability and expectation are denoted by $\mathbb{P}$ and $\mathbb{E}$, respectively. For any integer $K\in\mathbb{N}$, the set $\{1,2,3,\cdots,K\}$ is represented by $[K]$. The notation $\left\|\mathbf{x}\right\|_{q}$, $\left\|\mathbf{X}\right\|_{F}$, $\left\|\mathbf{X}\right\|_{op}$ represent the $\ell_{p}$ norm of a vector $\mathbf{x}$, Frobenius norm and operator norm of matrix $\mathbf{X}$ respectively. Here we use $\mathbf{1}(\cdot)$ to display the indicator function. For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}\lesssim b_{n}$ means $a_{n}\leq Cb_{n}$ for some constant $C$.
2 Related work and main contributions
@@
The threshold for exact recovery in the weighted/labeled SBM was first investigated by Jog and Loh (2015), which extends the results established in Zhang and Zhou (2016) for the SBM. Moreover, Yun and Proutiere (2016) derived the optimal misclassification rate for the LSBM and introduced a two-stage algorithm to achieve the information-theoretic limit. Besides, the Contextual-SBM, a generalization of SBM that incorporates additive node attributes to provide further information for identifying communities, has been explored in various studies Braun et al. (2022); Deshpande et al. (2018); Dreveton et al. (2024); Abbe et al. (2022). However, to the best of our knowledge, the optimal misclassification rate for models with general non-binary labeled edges combined with independent arbitrary node attributes still remains unexplored. Our Theorem 1 establishes that the lower bound for the expected number of misclassified nodes (as defined in Eq. 2) is $n\exp\left(-(1+\epsilon_{n})D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})n\right)$ for some sequence $\epsilon_{n}$ approaching zero. Here the key quantity $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ (defined in (4)) is the divergence defined based on the minimal value for the sum of two KL divergences, representing the topological and attribute information respectively. Furthermore, Remark 1 reveals a connection between the minimal value of sum of KL divergences over constrained sets and the Chernoff divergence, extending the results of (Yun and Proutiere, 2016, Claim $4$) to our settings and consequently prove that the aformentioned lower bound indeed match the optimal misclassification rate in the LSBM (Zhang and Zhou (2016); Yun and Proutiere (2016)) and the Gaussian mixture model (Lu and Zhou (2016)). Our main technical contribution stems in the analysis to the problem (4). It is worth noting that, while the optimal solution for $D_{A}$ has been investigated in Yun and Proutiere (2016), the impact of the newly introduced term $D_{X}$ on the main results remains unclear. The additional term introduces further challenges, as it involves in solving an optimization problem in function spaces, which is significantly more complex. In addition to establishing the theoretical lower bound, we also apply spectral-based methods to the proposed aggregated latent factor model, achieving a polynomial misclassification rate. Despite the suboptimality of the result, the output of the spectral-based method provides a valuable initialization for further refinement in a subsequent step (e.g., Gao et al. (2018); Yun and Proutiere (2016); Braun et al. (2022)).
@@ 3.1 Model definition
We largely follow the notation introduced in Yun and Proutiere (2016) to describe the LSBM. Denote $\mathcal{V}$ as the set of nodes constrained to $\textit{card}(\mathcal{V})=n$ and let $\{\mathcal{V}_{k}\}_{k\in[K]}$ represent the $K$ disjoint communities. Here each node is independently assigned to a community $\mathcal{V}_{k}$ with probability $\alpha_{k}$, where $k\in[K]$ and $\mathbf{1}^{T}\mathbf{\alpha}=1$, and this assignment is independent of the number of nodes $n$. Each edge $(i,j)\in\mathcal{V}^{2}$ is assigned to a label $l\in\{0,1,\cdots,L\}$, where the probability of labeling depends on $i$, $j$, and $l$, and is denoted by $\mathbf{P}(\sigma(i),\sigma(j),l)$. Here, $\mathbf{P}\in\mathbb{R}^{K\times K\times(L+1)}$, and $\sigma(\cdot)$ is the ground truth indicator function. We summarize the above in the definition of LSBM as follows:
Definition 1 (LSBM).
@@
Let $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$ and we denote $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}\sim\textit{LSBM}\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$ if $\forall l\in\{0,1,\cdots,L\}$ $\mathbf{A}_{l}\in\left\{0,1\right\}^{n\times n}$ is symmertical and its upper triangular elements are independent, such that:
$\mathbb{P}(\mathbf{A}_{l}(i,j)=1)=\mathbf{P}(\sigma(i),\sigma(j),l),~{}\sum_{l=0}^{L}\mathbf{A}_{l}(i,j)=1,$
$\mathbb{P}(\sigma(i)=k)=\alpha_{k},$ (1)
$\forall i\neq j\in[n],\,k\in[K]$. Here we assume $\mathbf{A}_{0}(i,i)=1$ for all $i\in[K]$ to eliminate the self-loops within the graph.
@@
In addition to topological information, contextual node attributes are also incorporated into the final model. We assume that each node $i$ is associated with an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the graph $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$ is considered independent of the sets of attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$ given the ground truth indicator function $\sigma(\cdot)$. This setting was initially proposed in Zhang et al. (2016) and has since been discussed in subsequent works Binkiewicz et al. (2017); Yan and Sarkar (2021); Weng and Feng (2016); Deshpande et al. (2018).
Definition 2 (CLSBM).
@@
Let $d\geq 1$, $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$, $\boldsymbol{\mu}\in\mathbb{R}^{d\times K}$, and we denote ($(\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}},\left\{\mathbf{x}_{i}\right\}_{i\in[n]})\sim\textit{CLSBM}(L,d,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot))$, if $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}\sim\textit{LSBM}\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$, and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\sim\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
Let $\bar{\sigma}(\cdot)$ be an assignment vector function that indicates the assignment for each node and is estimated from the observed edge labels $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}$ and node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$. In the next section, we derive an information-theoretic lower bound for the expected number of misclassified nodes, valid for all the possible choices of $\tilde{\sigma}(\cdot)$, defined as follows
@@
$\bar{s}:=\inf_{\begin{subarray}{c}\pi\in S_{K}\\ \tilde{\sigma}\in\mathcal{F}_{I}\end{subarray}}~{}\mathbb{E}\left[\sum_{i\in[n]}\mathbf{1}_{\{\pi(\tilde{\sigma}(i))\neq\sigma(i)\}}\right].$ (2)
@@
Here, $S_{K}$ represents the set of all permutations of $[K]$, and $\mathcal{F}_{I}$ denotes the set of all indicator functions. Before presenting the main results, we first outline our assumption as follows.
@@
$\forall l\in\{0,1,\cdots,L\},~{}\forall i,j,k\in[K],\quad\frac{\mathbf{P}(i,j,l)}{\mathbf{P}(i,k,l)}\leq\eta_{1},\quad\textit{and}\quad\left\|\boldsymbol{\mu}_{i}\right\|_{2}\leq\eta_{2}.$ (3)
Although counterintuitive, the aforementioned assumption implies that the signal contributed by any one of the $K$ communities cannot be excessively strong for both the network and node attributes. This assumption arises from the limitations of our proof techniques, and it could potentially be eliminated with improved proof mechanisms. Specifically, the assumption for $\boldsymbol{\mu}$ is newly introduced here to bound the second-order moments of the pseudo-likelihood estimator for the perturbed model in the proof of Theorem 1. Similarly, the assumption regarding $\mathbf{P}$ originates from Yun and Proutiere (2016), serves the the same purpose. It is also worth mentioning that similar assumptions on $\boldsymbol{\mu}$ are commonly found in previous literature, such as Abbe et al. (2022).
3.2 Lower bound on the number of misclassification nodes
In this section we provide one of our main results. We start with the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$:
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{k_{1}\neq k_{2}\in[K]}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}_{+}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}),$ (4)